OpenStudy (anonymous):

A brownie recipe asks for one and two thirds times as much sugar as chocolate chips. If three and three quarters cups of sugar is used, what quantity of chocolate chips would then be needed, according to the recipe?

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@Leonhart i tried, and it was wrong.

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

@animalain

jimthompson5910 (jim_thompson5910):

S = (1+2/3)*C S = (3/3+2/3)*C S = (5/3)*C 3 + 3/4 = (5/3)*C 12/4 + 3/4 = (5/3)*C 15/4 = (5/3)*C (15/4)*(3/5) = C C = 9/4 C = 2 + 1/4 C = 2 1/4 So you need 2 and 1/4 cups of chocolate chips